passing through (-1,3) and parallel to the line whose equation is 3x-2y-5=0
Answer by mananth(16946)

3 x -2 y = 5
Find the slope of this line
make y the subject
-2 y = -3 x + 5
Divide by -2
y = 3/2 x -5/2
Compare this equation with y=mx+b
slope m = 3/2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 3/2
m= 3/2 ,point ( -1 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = -3/ 2 + b
b= 4.5
m= 3/2
Plug value of the slope and b in y = mx +b
The required equation is y = 3/2 x + 9/ 2

m.ananth@hotmail.ca